OpenStudy (anonymous):

If f(x)=x² + 1, find f(a + 1). A. a² + 2 B. a²+a + 1 C. a²+2a + 2

OpenStudy (callisto):

f(a + 1) = (a + 1)² + 1 = a²+2a+1+1 = a²+2a + 2 =C

OpenStudy (anonymous):

C.